Starting from rest, how long does it take a moving man to reach a velocity of 5 m/sec if his acceleration is 1.75 m/sec^2?
1
Expert's answer
2010-10-02T22:59:55-0400
According to acceleration formula :

a = v - v0 / t - t0 ; according to our situation - t0 = 0; v0 = 0 so

a = v / t , so t = a/v = 1.75 / 5 = 0.35